in how many ways can we seat three of five boys and three of five girls if boys and girls must alternate?
Answer by stanbon(75887)

in how many ways can we seat three of five boys and three of five girls if boys and girls must alternate?
--------------
Pick a boy: 5 ways
Pick a girl: 5 ways
Pick a boy: 4 ways
Pick a girl: 4 ways
Pick a boy: 3 ways
Pick a girl: 3 ways
---------------------
# of ways = 5^2*4^2*3^2 = 25*16*9 = 3600
==========
Cheers,
Stan H.